### v1

@@ 4.2.1 Identifying useful Information in Bug Reports
While reproducing deep learning bugs, we kept track of information from bug reports that helped us reproduce them. After reproducing 85 bugs successfully, we have identified five useful pieces of information that can improve the chance of reproducing a bug. We discuss these factors in detail below.
Data (F): Data is one of the essential factors in ensuring the reproducibility of deep learning bugs. Deep learning systems heavily rely on the data breck2019data , and reproducing deep learning bugs becomes easier with access to the original data. Data helps us reproduce the deep learning bugs by providing the exact sample inputs that trigger the erroneous behaviour. By understanding the training data distributions and ranges, we can reconstruct the original training environment, which is crucial for reproducibility. However, the issue reports often lack direct information about the data. To address this problem, we collect information about the data by extracting the shape of the data, data distribution, type of variables and their corresponding ranges from the issue description. Leveraging this information and our proposed edit actions, namely Input Data Generation (A) and Dataset Procurement (A), we can generate or obtain the necessary data to reproduce deep learning bugs. According to our investigation, 73% of our reproduced bugs contained information about the data (see Table. 6) in their issue reports. For example, in the Stack Overflow post (Issue #43464835), the reporter has provided the dimensions and sample row of the training dataset, as shown below.
@@
Using this information and our proposed edit action Input Data Generation (A), we first generated a data frame of size (300, 5, 720) that contains values in the range from 0 to 400. Then, by leveraging the generated data and other useful information, we successfully reproduced the corresponding bug.
Model (F): The model architecture describes the components of a deep learning model and how they transform inputs into outputs. Understanding the model architecture is crucial for reproducing deep learning bugs, as it provides insights into the model’s components, connectivity, and the required neural network architecture. According to our investigation, 65% of our reproduced bugs contained information about the model architecture (see Table. 6) in their issue reports. However, the complete source code implementing the model architecture might not always be available in the issue reports. To overcome this limitation, we carefully gather information about a model’s architecture, i.e., the number of layers, layer properties and activation function from the issue description. Leveraging this information and our proposed edit action Neural Network Construction (A), we reconstruct the model and reproduce the deep learning bugs. For example, let us consider the following text from a Stack Overflow post (Issue #63204176) that mentions the use of multinomial logistic regression on the MNIST dataset, as shown below. Unfortunately, the reporter fails to provide the code snippet for the model.
@@
Despite the absence of the code snippet, the issue description provides useful hints about the model architecture and dataset. Using them and our edit action – Neural Network Construction (A), we successfully reproduced the bug.
Hyperparameters (F): Hyperparameters play a crucial role in controlling the learning process and model behaviour during training. They encompass parameters such as learning rate, batch size, number of epochs, optimizer, regularization techniques, and loss functions. The specific values chosen for these hyperparameters significantly impact a model’s performance, training process, and bug manifestation. Thus, reporting the complete set of hyperparameters is essential to help reproduce deep learning bugs. According to our investigation, 52% of our reproduced bugs contain information about the hyperparameters used in their issue reports. For the 48% of bug reports that did not include hyperparameter information, we used the Hyperparameter Initialization edit action (A) to reproduce the bugs by initializing the hyperparameters with default values. In the example Stack Overflow post (Issue #65993928), we can see how hyperparameters can play a crucial role in reproducing the bug.
@@
Code Snippet (F): Code snippets are critical for reproducing bugs, as highlighted by the fact that 98% of professional developers consider them an essential component of bug reproducibility soltani2020significance . They include the data preprocessing, data splitting technique, code used for training the model, and implementation of the evaluation metrics. However, from our manual bug reproduction, we observe that even though code snippets are present in 79% of the bug reports, only 9.41% of them can be used verbatim for bug reproduction. To address this limitation, we use several of our proposed edit actions, such as Import Addition and Dependency Resolution (A), Logging (A), Obsolete Parameter Removal (A), Compiler Error Resolution (A) and Version Migration (A). These edit actions help us fix the errors in the code and make the code compileable and runnable.
@@
Logs (F): Logs provide a real-time record of the model’s behaviour during training and inference. Traditional software logs consist of information such as event logs and stack traces, whereas deep learning systems consist of compiler error logs, training error logs, and evaluation logs Chen_Jiang_2022 . Sharing these logs is crucial for reproducing deep learning bugs as they allow us to verify if we can reproduce the same erroneous behaviour reported in the original issue. From our manual bug reproduction, we discover that 86% of our reproduced bugs contain the necessary logs in their issue reports. Such a high presence of logs in our dataset of reproduced bugs highlights the importance of logging in deep learning bug reproduction. By matching the logs from the original issue and the logs from reproduction on our local machines, we were able to confirm the presence of several bugs in the deep learning systems.
@@
| Prevalence | 72.9% | 64.7% | 52.7% | 78.8% | 85.8% |
4.2.2 Relationship between Useful Information and Type of Bugs
For training bugs, the data has a confidence value of 80.77% according to our generated rules (check Table. 7). This high confidence highlights the significance of data in reproducing the numerous training issues that can manifest during model development. The details about the data, such as the number of samples, class distribution, feature distributions, data splitting ratios, and preprocessing steps, are instrumental in reproducing training bugs. For example, training a classification model on a highly skewed dataset is prone to overfitting. With the knowledge of the data distribution, we can recreate a representative dataset, which can help us reproduce the bugs, even if the actual dataset is not available.
For tensor bugs, data characteristics have an even higher confidence of 95.24%, according to our generated rules (Table. 7). This suggests that tensor attributes like shape, data type, sparsity, value ranges and origin are pivotal for reliably reproducing many bugs stemming from invalid dimensions or precision issues. For instance, bugs arising from tensor shape mismatches can emerge if the shape of the input data does not match the expected input shape. Furthermore, real-world data is often more vulnerable to human error and biases compared to synthetic data talwar2020evaluating . Therefore, comprehensive documentation and availability of these salient data characteristics are imperative for the reliable reproduction of the numerous training and tensor bugs in deep learning systems.
Model: A neural network model’s architecture and implementation details are crucial to reproducing model and API bugs. The model architecture has 70% and 72.71% confidence values for model and API bugs, respectively, according to our generated rules. These high values signify that access to model details is vital for reliably reproducing issues stemming from model capacity, connectivity, and API usage.
Code Snippet: Code snippets are invaluable for reproducing deep learning bugs since they isolate and encapsulate the core logic that triggers them. Developers can demonstrate and share the essence of buggy behaviour by creating a minimal reproducible example. Code snippets have very high confidence values of 92.3% for training bugs, 66.67% for tensor bugs, 70.0% for model bugs, and 81.8% for API bugs according to our generated rules. These high values across all bug types signify that code snippets are critical for reliably reproducing deep learning bugs.
Logs: Logs play a fundamental role in deep learning by facilitating the reproduction and resolution of bugs. They comprehensively record every detail during model training, capturing information about various conditions, configurations, and events before the failures. Logs have high confidence values of 84.77% for training bugs, 85.71% for tensor bugs, 80% for model bugs, and 90.91% for API bugs according to our generated rules. Such high confidence levels across all bug types highlight that comprehensive log recording is important for the reliable reproduction of deep-learning bugs.
@@ 4.2.2 Relationship between Useful Information and Type of Bugs
| Input Data Generation (0.6154) | Neural Network Definition (0.75) |
@@
| Version Migration (0.3477) | Hyperparameter Initialization (0.50) |
| Hyperparameter Initialization (0.3077) | Logging (0.45) |
| Compiler Error Resolution (0.2692) | Dataset Procurement (0.40) |
| Tensor | API |
| Import Addition (0.4545) | Input Data Generation (0.5714) |
| Input Data Generation (0.3636) | Hyperparameter Initialization (0.5238) |
| Hyperparameter Initialization (0.3636) | Import Addition (0.4762) |
@@
| Logging (0.1818) | Obsolete Parameter Removal (0.3333) |
@@ 4.2.3 Relationship between Edit Actions and Type of Bugs
After determining the association between the useful information needed and the type of bug, we derived the relationship between the type of bug and the edit actions. We use the Apriori algorithm to determine the relationship, as done earlier. Table. 8 summarizes our findings, and we discuss them in detail below.
@@
Training Bug: Input data generation (61.54% confidence) is the most frequently used edit action for reproducing training bugs. However, our analysis (Table. 6) shows that 73% of these reports contain information about the data. Leveraging this information and our proposed edit action – Input Data Generation (A), we can reproduce the training bugs in deep learning systems. Furthermore, import addition (42.3% confidence) and version migration (34.7% confidence) are other edit actions which are often used to reproduce training bugs. Although 79% of bug reports contain code snippets, they are often incomplete. To reproduce the bug effectively, we thus need to complete these snippets by adding imports and migrating them to the latest versions of libraries and frameworks. Finally, hyperparameter initialization (30.77% confidence) and compiler error resolution (26.92% confidence) may also help reproduce training bugs by initializing the missing hyperparameters and resolving compile-time errors from the code snippet.
@@
Model Bug: Model bugs are often reproduced by modifying a neural network’s architecture, as shown by the high confidence value (75%) for the edit action – neural network construction. Since model bugs are primarily caused by errors in the neural network architecture taxonomyRealFaults , reconstructing the neural network might be the first step to reproduce them. Additionally, import addition (55% confidence) and initializing hyperparameters (50% confidence) are typical edit actions to reproduce the model bugs. Additionally, logging can also help the reproduction of model bugs by recording important details like layer outputs, gradients, and parameter values during training. These recorded values can then be compared between buggy and bug-free models, which can confirm the presence of a bug in the model. Thus, model bugs can be reliably reproduced through several edits that construct the neural network architecture, initialize hyperparameters, add logging, and procure datasets.
@@
Tensor Bug: Data is the most important information for reproducing tensor bugs, as shown by the strong correlation between Tensor Bug & Data (see Table. 7). This phenomenon can be attributed to the fact that tensor bugs primarily relate to the data taxonomyRealFaults . Since tensor bugs are so data-dependent, dataset procurement emerges as an important edit action to reproduce them. Bug reports reference the exact dataset used (if well-known) or describe the data type, shape and distribution. With this information, input data generation and dataset procurement can generate or procure the data required for the reproduction of the tensor bugs. Like other deep learning bugs, import addition (45.45% confidence) and hyperparameter initialization (36.36% confidence) also play a key role in completing the code snippet and reproducing the tensor bugs. Thus, tensor bugs can be reliably reproduced through several edits that procure the datasets, generate input data, add logging, import the required dependencies, and initialize the hyperparameters.
@@
API Bug: Since API Bugs stem from an improper usage of application programming interfaces (APIs), they do not focus as heavily on the model training process. As a result, bug reports might lack detailed information about the input data or hyperparameters used. This encourages the use of common edit actions such as input data generation (57.14% confidence) and hyperparameter initialization (52.38% confidence) when reproducing API bugs. Additionally, code snippets in bug reports are often incomplete, necessitating edit actions like import addition (47.62% confidence) and logging (42.86% confidence) to trace the intermediate states and outputs of the API. Thus, API bugs can be reliably reproduced through several edit actions that generate input data, import the required dependencies, add logging, initialize hyperparameters and remove obsolete parameters.
@@
4.3 RQ3: Do the suggested edit actions and information categories contribute to enhancing the reproducibility of deep learning bugs in practice?
@@
Table. 9 shows the bug reproducibility rate of the control and experimental groups across different sets of bugs. We found that developers in the control group could reproduce 72.91% bugs without hints. In contrast, the developers in the experimental group could reproduce 95.83% of bugs with our hints - a 22.92% increase. This significant increase in reproducibility rate demonstrates that our identified edit actions and information improved developers’ ability to reproduce deep learning bugs.
| % of bugs successfully reproduced by Control group | % of bugs successfully reproduced by Experimental group |
% Increase |
| Set 1 |
75.00 |
100.00 |
25.00 |
| Set 2 |
66.66 |
83.33 |
16.66 |
| Set 3 |
83.33 |
100.00 |
16.66 |
| Set 4 |
66.66 |
100.00 |
33.33 |
@@
We also calculate the time the control and experimental groups took to reproduce their assigned bugs. This helps us assess our findings’ benefits in reducing the time required for deep learning bug reproduction. Table. 10 shows the time taken to reproduce bugs by the control and experimental groups. The experimental group outpaced the control group in bug reproduction across all sets, with the most notable difference in Set 4, where the experimental group was 30.16% faster. On average, the control group lagged behind the experimental group by 24.35% across all sets. These results demonstrate that our recommended edit actions and information categories enabled the experimental group to reproduce deep learning bugs much faster than the control group that did not receive this information.

### v2

@@ 5.2.1 Identifying useful Information in Bug Reports
While reproducing deep learning bugs, we kept track of information from bug reports that helped us reproduce them. After reproducing 148 bugs successfully, we have identified five useful pieces of information that can improve the chance of reproducing a bug. We discuss these factors in detail below.
Data (F1): Data is one of the essential factors in ensuring the reproducibility of deep learning bugs. Deep learning systems heavily rely on the data breck2019data , and reproducing deep learning bugs becomes easier with access to the original data. Data helps us reproduce the deep learning bugs by providing the exact sample inputs that trigger the erroneous behaviour. By understanding the training data distributions and ranges, we can reconstruct the original training environment, which is crucial for reproducibility. However, the issue reports often lack direct information about the data. To address this problem, we collect information about the data by extracting the shape of the data, data distribution, type of variables and their corresponding ranges from the issue description. Leveraging this information and our proposed edit actions, namely Input Data Generation (A1) and Dataset Procurement (A8), we can generate or obtain the necessary data to reproduce deep learning bugs. According to our investigation, 77% of our reproduced bugs contained information about the data (see Table. 7) in their issue reports. For example, in the Stack Overflow post (Issue #43464835), the reporter has provided the dimensions and sample row of the training dataset, as shown below.
@@
Using this information and our proposed edit action Input Data Generation (A1), we first generated a data frame of size (300, 5, 720) that contains values in the range from 0 to 400. Then, by leveraging the generated data and other useful information, we successfully reproduced the corresponding bug.
Model (F2): The model architecture describes the components of a deep learning model and how they transform inputs into outputs. Understanding the model architecture is crucial for reproducing deep learning bugs, as it provides insights into the model’s components, connectivity, and the required neural network architecture. According to our investigation, 58% of our reproduced bugs contained information about the model architecture (see Table. 7) in their issue reports. However, the complete source code implementing the model architecture might not always be available in the issue reports. To overcome this limitation, we carefully gather information about a model’s architecture, i.e., the number of layers, layer properties and activation function from the issue description. Leveraging this information and our proposed edit action Neural Network Construction (A2), we reconstruct the model and reproduce the deep learning bugs. For example, let us consider the following text from a Stack Overflow post (Issue #63204176) that mentions the use of multinomial logistic regression on the MNIST dataset, as shown below. Unfortunately, the reporter fails to provide the code snippet for the model.
@@
Despite the absence of the code snippet, the issue description provides useful hints about the model architecture and dataset. Using them and our edit action – Neural Network Construction (A2), we successfully reproduced the bug.
Hyperparameters (F3): Hyperparameters play a crucial role in controlling the learning process and model behaviour during training. They encompass parameters such as learning rate, batch size, number of epochs, optimizer, regularization techniques, and loss functions. The specific values chosen for these hyperparameters significantly impact a model’s performance, training process, and bug manifestation. Thus, reporting the complete set of hyperparameters is essential to help reproduce deep learning bugs. According to our investigation, 48% of our reproduced bugs contain information about the hyperparameters used in their issue reports. For the 52% of bug reports that did not include hyperparameter information, we used the Hyperparameter Initialization edit action (A2) to reproduce the bugs by initializing the hyperparameters with default values. In the example Stack Overflow post (Issue #65993928), we can see how hyperparameters can play a crucial role in reproducing the bug.
@@
Code Snippet (F4): Code snippets are critical for reproducing bugs, as highlighted by the fact that 98% of professional developers consider them an essential component of bug reproducibility soltani2020significance . They include the data preprocessing, data splitting technique, code used for training the model, and implementation of the evaluation metrics. However, from our manual bug reproduction, we observe that even though code snippets are present in 82% of the bug reports, only 9.41% of them can be used verbatim for bug reproduction. To address this limitation, we use several of our proposed edit actions, such as Import Addition and Dependency Resolution (A4), Logging (A5), Obsolete Parameter Removal (A6), Compiler Error Resolution (A7) and Version Migration (A10). These edit actions help us fix the errors in the code and make the code compileable and runnable.
@@
Logs (F5): Logs provide a real-time record of the model’s behaviour during training and inference. Traditional software logs consist of information such as event logs and stack traces, whereas deep learning systems consist of compiler error logs, training error logs, and evaluation logs Chen_Jiang_2022 . Sharing these logs is crucial for reproducing deep learning bugs as they allow us to verify if we can reproduce the same erroneous behaviour reported in the original issue. From our manual bug reproduction, we discover that 88% of our reproduced bugs contain the necessary logs in their issue reports. Such a high presence of logs in our dataset of reproduced bugs highlights the importance of logging in deep learning bug reproduction. By matching the logs from the original issue and the logs from reproduction on our local machines, we were able to confirm the presence of several bugs in the deep learning systems.
@@
| Prevalence | 77.4% | 58.1% | 47.9% | 82.1% | 87.6% |
@@
5.2.2 Relationship between Useful Information and Type of Bugs
For training bugs, the data has a confidence value of 82.00% according to our generated rules (check Table. 8). This high confidence highlights the significance of data in reproducing the numerous training issues that can manifest during model development. The details about the data, such as the number of samples, class distribution, feature distributions, data splitting ratios, and preprocessing steps, are instrumental in reproducing training bugs. For example, training a classification model on a highly skewed dataset is prone to overfitting. With the knowledge of the data distribution, we can recreate a representative dataset, which can help us reproduce the bugs, even if the actual dataset is not available.
For tensor bugs, data characteristics have an even higher confidence of 96.55%, according to our generated rules (Table. 8). This suggests that tensor attributes like shape, data type, sparsity, value ranges and origin are pivotal for reliably reproducing many bugs stemming from invalid dimensions or precision issues. For instance, bugs arising from tensor shape mismatches can emerge if the shape of the input data does not match the expected input shape. Furthermore, real-world data is often more vulnerable to human error and biases compared to synthetic data talwar2020evaluating . Therefore, comprehensive documentation and availability of these salient data characteristics are imperative for the reliable reproduction of the numerous training and tensor bugs in deep learning systems.
Model: A neural network model’s architecture and implementation details are crucial to reproducing model and API bugs. The model architecture has 78.57% and 70.00% confidence values for model and API bugs, respectively, according to our generated rules. These high values signify that access to model details is vital for reliably reproducing issues stemming from model capacity, connectivity, and API usage.
Code Snippet: Code snippets are invaluable for reproducing deep learning bugs since they isolate and encapsulate the core logic that triggers them. Developers can demonstrate and share the essence of buggy behaviour by creating a minimal reproducible example. Code snippets have very high confidence values of 86.0% for training bugs, 72.41% for tensor bugs, 71.43% for model bugs, and 75.0% for API bugs, according to our generated rules. These high values across all bug types signify that code snippets are critical for reliably reproducing deep learning bugs.
Logs: Logs play a fundamental role in deep learning by facilitating the reproduction and resolution of bugs. They comprehensively record every detail during model training, capturing information about various conditions, configurations, and events before the failures. Logs have high confidence values of 76.00% for training bugs, 93.10% for tensor bugs, 78.57% for model bugs, and 85.00% for API bugs according to our generated rules. Such high confidence levels across all bug types highlight that comprehensive log recording is important for the reliable reproduction of deep-learning bugs.
@@ 5.2.2 Relationship between Useful Information and Type of Bugs
| Input Data Generation (0.5625) | Hyperparameter Initialization (0.5142) |
| Import Addition (0.4583) | Dataset Procurement (0.4390) |
| Compiler Error Resolution (0.3750) | Compiler Error Resolution (0.4146) |
| Dataset Procurement (0.3542) | Import Addition (0.4146) |
| Hyperparameter Initialization (0.3333) | Neural Network Construction (0.3659) |
| Tensor | API |
| Hyperparameter Initialization (0.5517) | Input Data Generation (0.40) |
| Input Data Generation (0.5172) | Hyperparameter Initialization (0.40) |
@@
| Obsolete Parameter Removal (0.3448) | Obsolete Parameter Removal (0.25) |
@@ 5.2.3 Relationship between Edit Actions and Type of Bugs
After determining association between the useful information and the type of bug, we derived the relationship between the edit actions and the type of bug. We use the Apriori algorithm to determine the relationship, as done earlier. Table. 9 summarizes our findings, and we discuss them in detail below.
@@
Training Bug: Input data generation (56.25% confidence) is the most frequently used edit action for reproducing training bugs. However, our analysis (Table. 7) shows that 77% of the bug reports/SO posts contain information about the data. Leveraging this information and our proposed edit action – Input Data Generation (A1), we can reproduce the training bugs in deep learning systems. Furthermore, import addition (45.8% confidence) and compiler error resolution (37.5% confidence) are other edit actions which are often used to reproduce training bugs. Although 79% of bug reports contain code snippets, they are often incomplete. To reproduce the bug effectively, we thus need to complete these snippets by adding the necessary import statements and migrating them to the latest versions of libraries and frameworks. Finally, dataset procurement (35.42% confidence) and hyperparameter initialization (33.33% confidence) may also help reproduce training bugs by procuring the dataset required and initializing the missing hyperparameters.
Since the training bugs have specific sub-faults, we manually analyse the Stack Overflow posts and Github issue reports for different sub-types and discuss the edit actions, which can be used to reproduce the specific sub-faults below.
Optimisation: Optimization bugs can be reproduced by employing a combination of edit actions, with a primary focus on three key areas: Hyperparameter Initialization (60.00% confidence), Input Data Generation (60.00% confidence), and Neural Network Construction (40.00% confidence). These actions involve configuring the optimizer and learning rate, creating representative training data, and building the model architecture, respectively. Additionally, logging (40.00% confidence) can help monitor the training process, while Import Addition (20.00% confidence), Compiler Error Resolution (20.00% confidence), and Version Migration (20.00% confidence) ensure compatibility within the environment. By systematically testing different optimizer configurations, generating appropriate input data, and constructing the relevant model architecture, developers can effectively reproduce optimization issues and gain insight into the underlying causes.
Loss Function: Loss function bugs, which can arise from incorrect loss calculations or suboptimal loss function choices, can be reproduced through a combination of edit actions. Hyperparameter Initialization (66.67% confidence) plays a crucial role in configuring the loss function while Logging (58.33% confidence) captures the output of intermediate loss values during training. Since the design of the output layer in the neural network directly influences the calculation of loss, Neural Network Construction (41.67% confidence) may be necessary to reproduce the bugs related to the loss function. Furthermore, it is essential to generate appropriate training data through Input Data Generation (33.33% confidence) to effectively reproduce bugs related to the loss function.
Hyperparameters: Hyperparameter sub-faults, such as suboptimal batch size, suboptimal number of epochs, and suboptimal learning rate, can be reproduced by utilizing the Hyperparameter Initialization edit action (64.29% confidence). This involves initializing various hyperparameters, such as batch size, epochs, and learning rate, that may be missing or suboptimal in the provided code. By experimenting with different commonly used values for these hyperparameters, we were able to reproduce them. In addition to that, Input Data Generation (50.00% confidence) proves to be valuable for reproducing hyperparameter bugs. Generating representative input data is crucial for triggering hyperparameters-related issues during the training process. Furthermore, actions like Import Addition (42.86% confidence), Logging (35.71% confidence), and Compiler Error Resolution (35.71% confidence) are helpful in ensuring that the code runs smoothly and allows for testing different hyperparameter configurations. These actions contribute to creating a compatible environment to address hyperparameter-related concerns effectively.
@@
Model Bug: Model bugs are often reproduced by using the edit action – hyperparameter initialization (51.4% confidence). From our manual analysis, we observed that the hyperparameters were not often reported for model bugs. Hence, hyperparameter initialization was often used to reproduce the model bugs. Additionally, dataset procurement (43.90% confidence) and compiler error resolution (41.46% confidence) are typical edit actions to reproduce the model bugs. Moreover, the code snippets for deep learning bugs are often incomplete, as observed in our manual analysis; hence, import addition (41.46% confidence) is a critical edit action for completing the code snippet and reproducing model bugs. Finally, model bugs might be reproduced by modifying a neural network’s architecture, as shown by the moderate confidence value (36.6%) for the edit action – neural network construction. Since model bugs are primarily caused by errors in the neural network architecture taxonomyRealFaults , reconstructing the neural network might be the first step to reproduce them. Thus, model bugs can be reliably reproduced through several edits that initialize hyperparameters, resolve compiler errors, procure datasets, add imports, and construct neural networks.
Similar to training bugs, model bugs also have specific sub-faults. Hence, we report the edit actions for reproducing specific sub-faults below.
Layer Type & Properties: To reproduce specific model bugs related to individual layers, such as incorrect layer types, suboptimal filter sizes, or inappropriate activation functions, it is necessary to carefully define the layers as part of the Neural Network Construction (78.57% confidence) action. By accurately specifying the layers in the model architecture, developers can trigger the desired layer-specific bugs for further analysis and resolution. It is also crucial to initialize the appropriate hyperparameters through the Hyperparameter Initialization (57.14% confidence) action. Additionally, obtaining the required input data shapes via Input Data Generation (42.86% confidence) is essential for triggering layer-specific issues during the training process. To ensure compatibility within the project environment, actions such as Compiler Error Resolution (42.86% confidence) and Version Migration (42.86% confidence) are beneficial for updating the layer definitions and reproducing any compatibility issues that may arise.
Model Type & Properties: Bugs associated with suboptimal model architecture, incorrect network structure, or missing layers can often be reproduced using the Neural Network Construction edit action (71.43% confidence). By carefully constructing the neural network based on the architecture details provided in a bug report, developers can reproduce sub-faults from the model bugs category. The interaction between the hyperparameters and the model architecture could be important for model bugs. Therefore, the Hyperparameter Initialization action (42.86% confidence) also plays a significant role in effectively reproducing model-related issues. To ensure compatibility with the latest library and framework version, actions such as Import Addition (35.71% confidence), Obsolete Parameter Removal (28.57% confidence), and Compiler Error Resolution (28.57% confidence) are essential. These actions help update the codebase and ensure that the necessary dependencies are met to define the desired model architecture accurately.
@@
Tensor Bug: Data is the most important information for reproducing tensor bugs, as shown by the strong correlation between Tensor Bug & Data (see Table. 8). This phenomenon can be attributed to the fact that tensor bugs primarily relate to the data taxonomyRealFaults . Since tensor bugs are so data-dependent, dataset procurement emerges as an important edit action to reproduce them. Bug reports reference the exact dataset used (if well-known) or describe the data type, shape and distribution. With this information, input data generation and dataset procurement can generate or procure the data required for the reproduction of the tensor bugs. Like other deep learning bugs, import addition (51.72% confidence) and hyperparameter initialization (55.17% confidence) also play a key role in completing the code snippet and reproducing the tensor bugs. Thus, tensor bugs can be reliably reproduced through several edits that procure the datasets, generate input data, add logging, import the required dependencies, and initialize the hyperparameters.
@@
API Bug: Since API Bugs stem from an improper usage of application programming interfaces (APIs), they do not focus as heavily on the model training process. As a result, bug reports might lack detailed information about the input data or hyperparameters used. This encourages the use of common edit actions such as input data generation (40.00% confidence) and hyperparameter initialization (40.00% confidence) when reproducing API bugs. Additionally, code snippets in bug reports are often incomplete, necessitating edit actions like import addition (35.00% confidence) and logging (25.00% confidence) to trace the intermediate states and outputs of the API. Thus, API bugs can be reliably reproduced through several edit actions that generate input data, import the required dependencies, add logging, initialize hyperparameters and remove obsolete parameters.
Summary of RQ2: By applying the Apriori algorithm on the data produced from our reproduction of 148 deep learning bugs, we identified the top 3 most important pieces of information and the top 5 edit actions needed to reproduce each category of deep learning bug. This provides insights into the missing information that should be solicited in bug reports as well as the edit actions required to reproduce specific bug types.
@@
5.3 RQ3: How do the suggested edit actions and component information affect the reproducibility of deep learning bugs?
@@
Table. 10 and Fig. 3 shows the bug reproducibility rate of the control and experimental groups across different sets of bugs. We found that developers in the control group could reproduce 72.91% bugs without hints. In contrast, the developers in the experimental group could reproduce 95.83% of bugs with our hints - a 22.92% increase. This significant increase in reproducibility rate demonstrates that our identified edit actions and information improved developers’ ability to reproduce deep learning bugs.
| % of bugs successfully reproduced by Control group | % of bugs successfully reproduced by Experimental group | % Increase | |
| Set 1 | 75.00 | 100.00 | 25.00 |
| Set 2 | 66.66 | 83.33 | 16.66 |
| Set 3 | 83.33 | 100.00 | 16.66 |
| Set 4 | 66.66 | 100.00 | 33.33 |
@@
Question: What are the challenges of bug reproduction in day-to-day activities?
R1: lack of debugging support and missing information
R2: data quality issues and lack of standardized debugging procedures and tools.
R3: no standardized debugging practices, and lack of clarity on the information needed, also lot of dependencies (libraries, framework, data, infra and so on)
R4: the flaky nature of deep learning models, and the unclear expectations of how model is supposed to behave.
R5: memory issues, documentation issues (missing information in issues), weak debugging support
R6: version control of deep learning is tricky, because of multiple snapshots of models, model management and reproducibility is tricky. also, the lack of standardized debugging practices makes it more tricky.
R7: distributed computing makes it difficult to find and reproduce the bugs. test coverage is also a problem as we cannot find the bugs properly because of lack of coverage and that is a problem with the reproduction of bugs.
@@
R1: I followed the guidance in the survey and used the hints to generate the appropriate training dataset, which then allowed me to reliably reproduce the problematic behavior during model training.
R2: had to add manual imports for torch and nn, and fixed the compiler errors related to imports, as highlighted by the hint provided
R3: Following the hint, I systematically generated all of the necessary input data that would be required in order to reliably reproduce the software bug during testing.
R4: The hint mentions the Iris dataset, so I used the edit action called “Dataset Procurement” and got the dataset, downloaded it and edited the code snippet to reproduce the bug.
R5: with the given hint, I generated the input data required for the bug reproduction.
R6: The imports were missing and as per the hint, data frame was generated for specific columns, which helped the resolution of the bug.
R7: As hinted in the survey, explicitly specifying the columns when constructing the data frame helped in the bug reproduction
We also calculate the time the control and experimental groups took to reproduce their assigned bugs. This helps us assess our findings’ benefits in reducing the time required for deep learning bug reproduction. Table. 11 and Fig. 4 show the time taken to reproduce bugs by the control and experimental groups. The experimental group outpaced the control group in bug reproduction across all sets, with the most notable difference in Set 4, where the experimental group was 30.16% faster. On average, the control group lagged behind the experimental group by 24.35% across all sets. These results demonstrate that our recommended edit actions and component information enabled the experimental group to reproduce deep learning bugs much faster than the control group that did not receive this information.